about triangle....
IF B IS TWICE AS LONG AS A AND B/C=4/5
WHAT IS THE PERIMETER OF THE TRIANGLE IF A=4 METERS?????????

Answer by Alan3354(21583)

IF B IS TWICE AS LONG AS A AND B/C=4/5
WHAT IS THE PERIMETER OF THE TRIANGLE IF A=4 METERS?
-------------
a = 4
b = 2a = 8
-----------
b/c = 4/5
5b = 4c
40 = 4c
c = 10
------
Perimeter = a+b+c = 22 meters